Naomi and Linda live 324 miles apart. They start at the same time and travel toward each other. Naomi's speed is 8 mph greater than Linda's. if they meet in 3 hours, find their speeds.
Found 2 solutions by jorel1380, Alan3354:
Answer by jorel1380(3719)

Let n be Linda's speed. Then Naomi's speed is n+8. So:
324/n+n+8=3
6n+24=324
6n=300
n=50
Linda is traveling at 50 mph; Naomi is traveling 58 mph. ☺☺☺☺

Answer by Alan3354(69443)

Naomi and Linda live 324 miles apart. They start at the same time and travel toward each other. Naomi's speed is 8 mph greater than Linda's. if they meet in 3 hours, find their speeds.
------------
Dividing miles by hours --> mi/hr
324/3 = 108 mi/hr
---
Moving in opposite directions --> speeds are added.
N + L = 108
N - L = 8
-------------------- Add